A car moving at 30 m/s uniformly slows down to a speed of 10 m/s in a time of 6 seconds. Determine the acceleration of the car.


1
Expert's answer
2021-10-20T16:39:58-0400
"a=\dfrac{v-v_0}{t}=\dfrac{10\ \dfrac{m}{s}-30\ \dfrac{m}{s}}{6\ s}=-3.3\ \dfrac{m}{s^2}."

The sign minus means that the car decelerates.